OpenStudy (anonymous):

solve for x 4|2x+5|-2=10

OpenStudy (anonymous):

@JuanitaM

OpenStudy (anonymous):

\[4\left| 2x+5 \right|=10+2 ,\left| 2x+5 \right|=\frac{ 12 }{ 4 }=3\] \[2x+5=\pm3\] solve these two equations and find the value of x

OpenStudy (anonymous):

2x = -2 x= -1

OpenStudy (anonymous):

2x+5=3 and 2x+5 -3 you get two solutions.

OpenStudy (anonymous):

one is correct. find the other also.

OpenStudy (anonymous):

2x+5=-3

OpenStudy (anonymous):

2x = -8 x = -4

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

alright thank you and one more question

OpenStudy (anonymous):

Can you check this for me?

OpenStudy (anonymous):

sure if i can.

OpenStudy (anonymous):

\[|10x + 5| \ge 25\]

OpenStudy (anonymous):

\[Solution: x \ge 2 \]

OpenStudy (anonymous):

\[\left| x \right|\ge a means x \ge a and x \le -a ,solve these two .\]

OpenStudy (anonymous):

\[ x \le -3 \]

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you are welcome.